Can you help me solve x/2+1/15=1/30X
Answer by checkley71(8403)

X/2+1/15=1/30X OR (15X+2)/30=1/30X OR 30=450X~2+60x or 450x~2+60x-30=0 or dividing all terms by 10 we get 45x~2+6x-3=0 or (5x-1)(9x+3)=0 or
5x=1 or x=1/5 & 9x=-3 x=-3/9 or x=-1/3
proof
(1/5)/2+1/15=1/(30x1/5) or 1/10+1/15=1/6 or 25/150=1/6 or 1/6=1/6 &
(-1/3)/2+1/15=1/(30x-1/3) or -1/6+1/15=1/-10 or -5/30+2/30=-1/10 or -3/30=-1/10
or -1/10=-1/10